OpenStudy (phagey):

The local grocery store offers the following products at the given prices. Find the more economical purchase and the unit cost. Barilla® Pasta, $1.69 for 12 ounces, or Cream Field® Pasta, $2.75 for 24 ounces Cream Field Pasta unit price = $.14 per ounce Barilla Pasta unit price = $.14 per ounce Cream Field Pasta unit price = $.11 per ounce Barilla Pasta unit price = $.11 per ounce

OpenStudy (anonymous):

I don't know. I'm not good at this stuff. sorry

OpenStudy (anonymous):

It could either be Cream Field Pasta unit price = $.11 per ounce or Barilla Pasta unit price = $.11 per ounce

OpenStudy (vickyisthesmartone123):

32 ounces for $2.99 unit cost = $0.09 per ounce- is the more economical purchase and the unit cost. ($2.99/32oz = $0.09/oz, $0.85/6oz = $0.14/oz) 32 oz for $2.99 = $0.09/oz is cheaper